OpenStudy (anonymous):

A recipe calls for 7/8 cup of sugar, and the baker has 2/3 cup of sugar. How much more sugar does the baker need to make the recipe? A) 5/8 c B) 3/5 c C) 5/24 c D) 1/4 c

ganeshie8 (ganeshie8):

hint- to compare fractions, convert them to same base

ganeshie8 (ganeshie8):

1) amount of sugar required : \(\frac{7}{8} = \frac{7*3}{8*3} = \frac{21}{24}\) 2) amount of sugar baker has : \(\frac{2}{3} = \frac{2*8}{3*8} = \frac{16}{24}\)

ganeshie8 (ganeshie8):

now you can easily see that the baker has only 16/24 cups of sugar, but the recipe needs 21/24 cups. so how much he needs more ?

ganeshie8 (ganeshie8):

still need help ?

OpenStudy (anonymous):

thanks!

ganeshie8 (ganeshie8):

np did you figure out.. hw much u got :P